Determine the point estimate of the population mean and margin of error for the confidence interval with lower bound: 19 and upper bound:27.

If you could please explain how to get answer :) , thank you!
Answer #1

Confidence interval = (19 , 27)

Point estimate = (19 + 27) / 2 = 23

Margin of error = (27 - 19) / 2 = 4